Pure benzene, C6H6, has a molar mass of 78.114 g mol-1, a density of 0.8765 g mL-1, a freezing point of 5.45°C, and a boiling point of 80.2°C. Its freezing point depression and boiling point elevation constants are: Kf = 5.07°C m-1; Kb = 2.53oC m-1. A solution was made by taking 33.88 g of an unknown nonelectrolyte and dissolving it in 175.0 g of benzene. The measured freezing point of the solution was 1.65oC. Calculate the molar mass of the unknown substance.
Answer #1

depression in freezing point, delta Tf = (5.45-1.65)oC

= 3.8 oC

delta Tf = kf*molality

3.8 = 5.07*m

m=0.7495

m = number of moles of solute / mass of solvent in kg

0.7495 = (mass of solute) / (molar mass of solute * mass of solvent in kg)

0.7495 = 33.88/(MM*0.175)

MM = 258.3 g/mol

Answer: 258.3 g/mol